OpenStudy (anonymous):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $40 million after 4 years, while profits increased $30 million per year. Select the correct graph of this function. There is a graph of a line that has a y intercept that is approximately $80 million and passes through the point of approximately x = 2 years and y = $140 million. There is a graph of a line that has a y intercept that is approximately negative $80 million and an x intercept that is approxim

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (mathmath333):

this can help http://openstudy.com/study#/updates/52945196e4b0999412bd7562

OpenStudy (anonymous):

thank you so much @mathmath333

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 45 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 3x + 45 f(x) = 3x + 45 f(x) = - 3x + 60 f(x) = 3x + 60

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). y - 4 = -one half(x + 6) y - 4 = 2(x + 6) y + 6 = -one half(x - 4) y + 6 = 2(x - 4)

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $80 million after 4 years, while profits increased $60 million per year. Select the correct graph of this function. There is a graph of a line that has a y intercept that is approximately negative $160 million and an x intercept that is approximately 2.7 years. There is a graph of a line that has a y intercept of $0 and an x intercept that is 0 years. The line passes through the point x = 2.5 and y = $50 million. There is a graph of a line that has a y intercept that is approximately negative $40 million and an x intercept that is approximately 1.3 years. There is a graph of a line that has a y intercept that is approximately $120 million and passes through the point x = 2 years and y = $100 million.